OpenStudy (anonymous):

For the Experiment, flip a coin until heads shows, assume that the probability of heads on one flip is 3/5. We define a RV (random variable) X= the number of flip. Find the probability that X will be no more than 10. You may find it helpful to use the formula fo a partial sum of a geometry series, or you may us another method ( but you must explain or show you computed your result, as in all of these problems).

OpenStudy (anonymous):

i guess you could compute the probability that it is greater than ten and subtract that from 1 not sure what is easiest though

OpenStudy (anonymous):

in other words, without using the partial sum of a geometric sequence, not sure how you would do it

OpenStudy (anonymous):

so i need the partial sum?

OpenStudy (anonymous):

damn site keeps crashing on me

OpenStudy (anonymous):

what i was trying to say was that it is worded in such a way that it suggests and easier method but i cannot think of one off the top of my head

OpenStudy (anonymous):

ok i see

OpenStudy (anonymous):

should be easy enough to add though btw did you get the other one?

OpenStudy (anonymous):

the other one no lol

OpenStudy (anonymous):

still thinking about it

OpenStudy (anonymous):

we can do that if you like this one should be very close to one

OpenStudy (anonymous):

\[\frac{3}{5}\times\sum_{k=0}^{9}\left(\frac{2}{5}\right)^k\] for this one

OpenStudy (anonymous):

the first one is actually easier it is \[\frac{3}{5}\sum_{k=0}^{\infty}\left(\frac{2}{5}\right)^k\]

OpenStudy (anonymous):

k i see

OpenStudy (anonymous):

if it is not clear how i got those, let me know

OpenStudy (anonymous):

k=9 and infinity ?

OpenStudy (anonymous):

the second one was for your previous problem the 9 is for this problem

OpenStudy (anonymous):

you mean the previous previous problem about prove ?

OpenStudy (anonymous):

right that one

OpenStudy (anonymous):

you had to prove that it added to one if i remember correctly

OpenStudy (anonymous):

yes to prove that all probability is 1using the same geometry series

OpenStudy (anonymous):

yes and you can sum \[\sum_{k=0}^{\infty}(\frac{2}{5})^k\] in your head

OpenStudy (anonymous):

you know how to do that?

OpenStudy (anonymous):

or more to the point, is it clear how to get the thing you need to sum?

OpenStudy (anonymous):

for this problem or the previous?

OpenStudy (anonymous):

either they are almost identical

OpenStudy (anonymous):

i guess i need to \frac{ 3 }{ 5 } sum_{k=0}^{9}( frac{ 2 }{ 5 })^(9

OpenStudy (anonymous):

\[\frac{ 3 }{ 5 } \sum_{k=0}^{9} (\frac{ 2 }{ 5 })^9\]

OpenStudy (anonymous):

yeah just asking if it is clear why that is the right thing it is easy enough to sum

OpenStudy (anonymous):

k so is (2/5)^9- (2/5)^0?

OpenStudy (anonymous):

the sum will be \[\frac{1-\left(\frac{2}{5}\right)^{10}}{1-\frac{2}{5}}\] then multiply by \(\frac{3}{5}\) that sum is very close to \(\frac{5}{3}\) so your answer will be almost 1

OpenStudy (anonymous):

since \((\frac{2}{5})^{10}\) is negligible

OpenStudy (anonymous):

i see as a large numbers

OpenStudy (anonymous):

not sure what you mean by large it is very close to \(\frac{5}{3}\) http://www.wolframalpha.com/input/?i=%281-%282%2F5%29^%2810%29%29%2F%283%2F5%29

OpenStudy (anonymous):

\[(2/5)^9\]

OpenStudy (anonymous):

that number is almost zero in but in fact it should be \[\frac{3}{5}\times\sum_{k=0}^{9}\left(\frac{2}{5}\right)^k=\frac{3}{5}\times \frac{1-\left(\frac{2}{5}\right)^{10}}{1-\frac{2}{5}}\]

OpenStudy (anonymous):

is 3/5* (2/5)^9

OpenStudy (anonymous):

?

OpenStudy (anonymous):

no it is the one i wrote

OpenStudy (anonymous):

first try prob is \(.6\) second try it is \(.4\times .6\) third try it is \(.4^2\times .6\) fourth try it is \(.4^3\times .6\) tenth try it is \(.4^9\times .6\) the sum is \[\frac{3}{5}\times\sum_{k=0}^{9}\left(\frac{2}{5}\right)^k=\frac{3}{5}\times \frac{1-\left(\frac{2}{5}\right)^{10}}{1-\frac{2}{5}}\] \[

OpenStudy (anonymous):

ok i see

OpenStudy (anonymous):

k good and as i said, that number is so close to one as to be essentially 1 which makes sense since if you flip a coin it is almost certain you will get a head by the tenth flip even more so with this bias coin that favors heads

OpenStudy (anonymous):

second .24, .96 and the .4^2 is larger numbers

OpenStudy (anonymous):

is like 1.5728...

OpenStudy (anonymous):

i mean the last one .4^9*.6